Two parallel and opposite forces, each 4,000 N are applied tangentially to the upper
and lower faces of a cubical metal block 25 cm on a side. Find the angle of shear and
the displacement of the upper surface relative to the lower surface. Shear modulus for
the metal is 80 GPa.
1
Expert's answer
2012-11-20T09:15:00-0500
Tan Q =(25×10 ̄ 7/ 25× 10 ̄ 2) = 8.0 × 10 ̄ 7 rad.
Displament = 4000 / (2.8 × 10 ̄ ²)²
80 Gpa = 64000 / Ʃ
Ʃ = 64000 / 80 G
= 80 × 10 ̄ 7 (25 × 10 ̄ ² )
= 2 × 10 ̄ 7.